OpenStudy (anonymous):

if you roll two dice, what is the probability that their sum will be 7(enter the answer as a fraction in simplest terms.)

OpenStudy (alfie):

7= 1+6 7= 6+1 7= 5+2 7= 2+5 7= 4+3 7= 3+4 So you can get seven in six different ways, which would be : \[\frac {1} {36} + \frac {1} {36} + \frac {1} {36} + \frac {1} {36} + \frac {1} {36} + \frac {1} {36} = \frac {6} {36} \] = \[\frac {1} {6}\]